### v1

@@ 2.3.2 Hierarchical Coordinate (HiCOO)
The hierarchical COO (HiCOO) format [18, 28] is a block-based sparse tensor representation, which employs multi-dimensional tiling schemes for data compression. Like COO, HiCOO is mode-agnostic but its compression efficacy depends completely on the properties of the target tensor, such as its shape, density, and data distribution, and determining the optimal parameters for compression (e.g., the tile size) is non-trivial. In some cases, rearranging the nonzero elements to create dense tiles is necessary to achieve any compression [28]. In addition, scheduling the resulting HiCOO blocks can suffer from limited parallelism, due to conflicting updates across blocks, as well as workload imbalance if some blocks have significantly more nonzero elements than others. Figure 3(b) shows the example sparse tensor encoded in the HiCOO format. The memory required to keep the hierarchical indices (i.e., , , etc.) can be lower than storing the actual indices (i.e., , , and ) only if each tile has a sufficient number of nonzero elements.
@@ 2.3.3 Compressed Sparse Fiber (CSF)
CSF stores a tensor as a collection of sub-trees, where each sub-tree represents a group of all nonzero elements that update the same factor matrix row. Given a CSF representation with a mode ordering of 1-2-3, where 1 is the root mode and 3 is the leaf mode, the root nodes represent the rows that will be updated, and the leaf nodes represent the nonzero elements that contribute to that update. Thus, iterating over the nonzero elements involves a bottom-up traversal of each sub-tree, such that at each non-leaf node, the partial results from its children are merged and pushed up, and this propagates until results from every node in the tree are merged at the root. Figure 3(c) shows illustrates the CSF sub-trees created from the example sparse tensor.
One advantage of CSF is that updates to the target factor matrix can be done asynchronously without requiring any synchronization by assigning one thread to each sub-tree. However, CSF requires copies of the tensor to maintain synchronization-free updates across every mode, which can be impractical for large tensors and/or devices with limited memory capacity (e.g., GPUs). Alternatively, the sub-trees can be traversed both bottom-up and top-down, meeting and merging at the tree level with the destination mode. While this approach allows a single copy of CSF (with the root mode chosen arbitrarily) to be used across all modes, it requires synchronization to avoid update conflicts and separate tree traversal algorithms. Additionally, regardless of the strategy used, CSF suffers from workload imbalance, as some sub-trees can have significantly more nonzero elements than others, and a thread assigned to a larger sub-tree will take longer to complete and keep other threads waiting.
@@ 3.1 ALTO Tensor Generation
Figure 4 shows the ALTO format for the sparse tensor from Figure 3(a). The multi-dimensional indices (, , and ) are color coded and the bit of their binary representation is denoted . Specifically, ALTO keeps the value of a nonzero element along with a linearized index, where each bit of this index is selected to partition the multi-dimensional space into two hyperplanes. For example, the ALTO encoding in Figure 4 uses a compact line of length (i.e., a -bit linearized index) to represent the target tensor of size . This index consists of three groups of bits with variable sizes (resolutions) to efficiently handle high-order data of arbitrary dimensionality. Within each bit group, ALTO arranges the modes in increasing order of their length (i.e., the shortest mode first), which is equivalent to partitioning the multi-dimensional space along the longest mode first. Such an encoding aims to generate a balanced linearization of irregular Cartesian spaces, where the position resolution of a nonzero element increases with every consecutive bit, starting from the most significant bit. Therefore, the line segments encode subspaces with mode intervals of equivalent lengths, e.g., the line segments , , and encode subspaces of , , and dimensions, respectively.
Due to this adaptive encoding of the multi-dimensional data, ALTO represents the resulting linearized index using the minimum number of bits, and it improves data locality across all modes of a given sparse tensor. Hence, a mode- tensor, whose dimensions are , can be efficiently represented using a single ALTO format with indexing metadata of size:
@@
where is the word size in bits. For example, on an architecture with byte-level addressing (i.e., bits), the sparse tensor in Figure 4 requires three bytes to store the mode indices (coordinates) for each nonzero element in the list-based COO format, whereas only a single byte is required to store the linearized index in the ALTO format: the metadata compression ratio of ALTO compared to the list-based COO is three.
@@ 4.1 Workload Partitioning and Scheduling
Once the linearized sparse tensor is divided into multiple line segments, ALTO identifies the basis mode intervals (coordinate ranges) of the multi-dimensional subspaces that correspond to these segments. For example, the line segments and correspond to three-dimensional subspaces bounded by the mode intervals and , respectively. While the derived multi-dimensional subspaces of the line segments may overlap, as highlighted in yellow in Figure 7, each nonzero element is assigned to exactly one line segment. That is, ALTO imposes a partitioning on a given linearized tensor that generates a disjoint set of non-overlapping and balanced line segments, yet it does not guarantee that such a partitioning will decompose the multi-dimensional space of the tensor into non-overlapping subspaces. In contrast, the prior sparse tensor formats decompose the multi-dimensional space into non-overlapping (yet highly imbalanced) regions, namely, tensor slices and fibers in CSF-based formats and multi-dimensional spatial blocks in block-based formats (e.g., HiCOO).
Formally, a set of line segments partitions an ALTO tensor , which encodes -dimensional sparse data, such that and , where . Each line segment is an ordered set of nonzero elements that are bounded in an -dimensional space by a set of closed mode intervals , where each mode interval is delineated by a start and an end . The intersection of two mode interval sets represents the subspace overlap between their corresponding line segments (partitions), as highlighted in yellow in Figure 7. This overlap information is used to more efficiently resolve conflicts between partitions, as described in 4.2.
@@ 4.2 Adaptive Conflict Resolution
Because processing the nonzero elements of a tensor in parallel (e.g., line 1 in Algorithm 3) can result in write conflicts across threads (e.g., line 3 in Algorithm 3), we devise an effective parallel execution and synchronization algorithm to handle these conflicts by exploiting the inherent data reuse of target tensors. Specifically, our adaptive algorithm chooses between 1) recursively traversing the tensor in ALTO order to maximize the reuse of both the input and output fibers, at the cost of global parallel reduction, or 2) traversing the tensor elements in output-oriented ordering, and only synchronizing at partition boundaries.
Algorithm 4 describes the proposed parallel execution scheme using a representative MTTKRP operation that works on a sparse tensor stored in the ALTO format. After ALTO imposes a partitioning on a given sparse tensor, as detailed in 4.1, each partition can be assigned to a different thread. To resolve the update/write conflicts that may happen during parallel sparse tensor computations, ALTO uses an adaptive conflict resolution approach that automatically selects the appropriate tensor traversal and synchronization technique (highlighted by the different gray backgrounds) across threads based on the reuse of the target fibers. This metric represents the average number of nonzero elements per fiber (the generalization of a matrix row/column) and it is estimated by simply dividing the total number of nonzero elements by the number of fibers along the target mode.
@@
ALTO considers the fiber reuse large enough to use local staging memory for conflict resolution, if the average number of nonzero elements per fiber is more than the maximum cost of using this two-stage (buffered) accumulation process, which consists of initialization (omitted for brevity), local accumulation (line 6 in Algorithm 4), and global accumulation (lines 14–18). In the worst (no reuse) case, the buffered accumulation cost is four memory operations (two read and two write operations). As explained in 4.1, each line segment is bounded in an -dimensional space by a set of closed mode intervals , which is computed during the partitioning of an ALTO tensor; thus, the size of the temporary storage accessed during the accumulation of ’s updates along a mode is directly determined by the mode interval (see lines 6 and 15).
@@ 4.3 Adaptive Memory Management
In many tensor decomposition algorithms, such as CP-APR, the intermediate results of tensor kernels are typically stored and then reused during the iterative optimization loop (as shown in Algorithm 2). However, storing these important calculations can substantially increase memory traffic and require prohibitive amount of memory, especially for large tensors and high decomposition ranks. Hence, we introduce an ALTO-based algorithm variant that recompute these values on-the-fly (ALTO-OTF). This in contrast to the traditional algorithm that pre-computes and reuses the intermediate values (ALTO-PRE). Moreover, we propose a heuristic to dynamically decide wither to reuse or recompute the intermediate results of tensor kernels based on the characteristics of the target sparse tensor data sets and computations.
To demonstrate our adaptive memory management technique, Algorithm 5 shows how the model update () kernel in CP-APR (Line 10 from Algorithm 2) is parallelized using the ALTO format. The matrix from Line 8 in Algorithm 2 calculates a dense matrix for a given mode that is the Khatri-Rao product (KRP) between all factor matrices excluding the mode- factor matrix. However, not every row of is required for a sparse tensor—only rows that correspond to nonzero elements are necessary and are actually calculated, leading to a matrix, where and are the number of nonzero elements and the decomposition rank, respectively. In Algorithm 5, if we select to use pre-computed , the kernel reads in the matrix row corresponding to the nonzero element from memory (Line 7); otherwise, it computes the required KRP from the factor matrices (Line 9) using the delineariezd coordinates. Pre-computing the matrix is simple; in line 8 of Algorithm 2 each nonzero element can calculate its Khatri-Rao product in parallel, using the equation from Line 9 in Algorithm 5.
Next, for each nonzero, the corresponding KRP row is used to update the matrix, and the conflicting updates are resolved using our adaptive conflict resolution as detailed in 4.2. Specifically, if there is high fiber reuse, recursive tensor traversal is conducted and the update is made to the temporary scratchpad memory Temp (Line 11), which are later reduced (Lines 20 to 24) to decrease memory contention; otherwise, output-oriented traversal is used to avoid synchronization and atomic operations are utilized to update the matrix (Line 13) only at the boundaries between different ALTO partitions/threads.
@@ 5 Evaluation
We evaluate ALTO-based tensor algorithms against the state-of-the-art sparse tensor libraries and representations in terms of parallel performance, tensor storage, and format generation cost. We conduct a thorough study of key tensor decomposition operations (2.2) and demonstrate the performance characteristics across the third and fourth generation of Intel Xeon Scalable processors, codenamed Ice Lake (ICX) and Sapphire Rapids (SPR) respectively.
Tensor | Dimensions | #NNZs | Density | Fib. reuse |
lbnl |
Limited |
nips |
High |
uber |
High |
chicago |
High |
vast |
High |
darpa |
Limited |
enron |
High |
lanl-2 |
High |
nell-2 |
High |
fb-m |
Limited |
flickr |
Limited |
deli |
Medium |
nell-1 |
Medium |
amazon |
High |
patents |
High |
High |
@@ 5.1.1 Platform
The experiments were conducted on an Intel Xeon Platinum 8360Y CPU, with Ice Lake (ICX) micro-architecture, and an Intel Xeon Platinum 8470 CPU, with Sapphire Rapids (SPR) micro-architecture. The ICX system has 256 GiB main memory and it consists of two sockets, each with a 54 MiB L3 cache and 36 physical cores running at a fixed frequency of 2.4 GHz for accurate measurements. The SPR system also comprises two sockets with 52 physical cores each, and its frequency was fixed to 2.4 GHz. All cores within a socket share a 105 MiB L3 cache and the overall main memory in the SPR server is 1 TB. Unless otherwise stated, the experiments use all hardware threads ( and , respectively) on the target platforms. Both servers run AlmaLinux 8.8 Linux distribution and they are configured to enable “Transparent Huge Pages” and to support two and four NUMA domains per socket on ICX and SPR, respectively.
The code is built using Intel C/C++ compiler (v2021.6.0) with the optimization flags -O3 -qopt-zmm-usage=high -xHost to fully utilize vector units. For performance counter measurements and thread pinning, we use the LIKWID tool suite v5.3 [43] and its SPR development branch.222https://github.com/RRZE-HPC/likwid/pull/524
@@ 5.1.3 Configurations
We evaluate the proposed ALTO format compared to the mode-agnostic COO and HiCOO formats [18, 22] as well as the mode-specific CSF formats [25, 20]. Specifically, we use the latest code of the state-of-the-art sparse tensor libraries for CPUs, namely, ParTI333Available at: https://github.com/hpcgarage/ParTI and SPLATT444Available at: https://github.com/ShadenSmith/splatt for normally distributed data and SparTen555Available at: https://github.com/sandialabs/sparten for non-negative count data. On the ICX and SPR systems, we evaluate the target data sets that can fit in memory for all tensor libraries. We report the best-achieved performance across the different configurations of the COO format; that is, with or without thread privatization (which keeps local copies of the output factor matrix). For the HiCOO format, its performance and storage are highly sensitive to the block and superblock (SB) sizes, which benefit from tuning. Since the current HiCOO implementation does not auto-tune these parameters, we use a block size of 128 () and two superblock sizes of (“HiCOO-SB10”) and (“HiCOO-SB14”) according to prior work [19]. We evaluate two variants of the mode-specific formats: CSF and CSF with tensor tilling (“CSF-tile”), both of which use representations (“SPLATT-ALL”) for an order- sparse tensor to achieve the best performance. Similar to previous studies [25, 20, 45], the experiments use double-precision arithmetic and -bit (native word) integers. To compute the CP-APR model for non-negative count data, we use -bit integers to represent the input tensor values. While the target data sets require a linearized index of size between and bits, we configured ALTO to select the size of its linearized index to be multiples of the native word size (i.e., and bits) for simplicity. We use a decomposition rank for all experiments and set the maximum number of inner iterations () in CP-APR to , as per prior work [14].

### v2

@@ 2.3.2 Hierarchical Coordinate (HiCOO)
HiCOO [15, 27] is a block-based sparse tensor format that employs multi-dimensional tiling for data compression. Like COO, HiCOO is mode-agnostic but its compression efficacy depends completely on the properties of the target tensor, such as its shape, density, and data distribution, and determining the optimal parameters for compression (e.g., the tile size) is non-trivial. In some cases, rearranging the nonzero elements to create dense tiles is necessary to achieve any compression [27]. In addition, scheduling the resulting HiCOO blocks can suffer from limited parallelism, due to conflicting updates across blocks, as well as workload imbalance if some blocks have significantly more nonzero elements than others. Figure 3(b) shows the example sparse tensor encoded in the HiCOO format. The memory required to keep the hierarchical indices (i.e., , , etc.) can be lower than storing the actual indices (i.e., , , and ) only if each tile has a sufficient number of nonzero elements.
@@ 2.3.3 Compressed Sparse Fiber (CSF)
CSF stores a tensor as a collection of sub-trees, where each sub-tree represents a group of all nonzero elements that update the same factor matrix row. Given a CSF representation with a mode ordering of 1-2-3, where 1 is the root mode and 3 is the leaf mode, the root nodes represent the rows that will be updated, and the leaf nodes represent the nonzero elements that contribute to that update. Thus, iterating over the nonzero elements involves a bottom-up traversal of each sub-tree, such that at each non-leaf node, the partial results from its children are merged and pushed up, and this propagates until results from every node in the tree are merged at the root. Figure 3(c) illustrates the CSF sub-trees created from the example sparse tensor.
One advantage of CSF is that updates to factor matrices can be done asynchronously by assigning one thread to each sub-tree. However, CSF requires tensor copies to maintain synchronization-free updates across every mode, which can be impractical for large tensors and/or devices with limited memory capacity (e.g., GPUs). Alternatively, the sub-trees can be traversed both bottom-up and top-down, merging partial results at the tree level corresponding to the destination mode. While this approach allows a single tensor copy (with the root mode chosen arbitrarily) to be used across all modes, it requires synchronization to avoid update conflicts and entirely different tree traversal algorithms [25]. Additionally, regardless of the strategy used, CSF suffers from workload imbalance, as some sub-trees can have significantly more nonzero elements than the others.
@@ 3.1 ALTO Tensor Generation
Figure 4 shows the ALTO format for the sparse tensor from Figure 3(a). The multi-dimensional indices (, , and ) are color coded and the bit of their binary representation is denoted . Specifically, ALTO keeps the value of a nonzero element along with a linearized index, where each bit of this index is selected to partition the multi-dimensional space into two hyperplanes. For example, the ALTO encoding in Figure 4 uses a compact line of length (i.e., a -bit linearized index) to represent the target tensor of size . This index consists of three groups of bits with variable sizes (resolutions) to efficiently handle high-order data of arbitrary dimensionality. Within each bit group, ALTO arranges the modes in an increasing order of their length (i.e., the shortest mode first), which is equivalent to partitioning the multi-dimensional space along the longest mode first. Such an encoding aims to generate a balanced linearization of irregular Cartesian spaces, where the position resolution of a nonzero element increases with every consecutive bit, starting from the most significant bit. Therefore, the line segments encode subspaces with mode intervals of equivalent lengths, e.g., the line segments , , and encode subspaces of , , and dimensions, respectively.
Due to this adaptive encoding, ALTO represents the resulting linearized index using the minimum number of bits, and it improves data locality across all modes of a given sparse tensor. Hence, a mode- tensor, whose dimensions are , can be efficiently represented using a single ALTO format with indexing metadata of size:
@@
where is the word size in bits. For example, on an architecture with byte-level addressing (i.e., bits), the sparse tensor in Figure 4 needs three bytes to store the mode indices (coordinates) for each nonzero element in the COO format, whereas only a single byte is required to store the linearized index in the ALTO format: the metadata compression ratio of ALTO compared to COO is three.
@@ 4.1 Workload Partitioning and Scheduling
Once the linearized sparse tensor is divided into multiple line segments, ALTO identifies the basis mode intervals (coordinate ranges) of the multi-dimensional subspaces that correspond to these segments. For example, the line segments and correspond to three-dimensional subspaces bounded by the mode intervals and , respectively. While the derived multi-dimensional subspaces of the line segments may overlap, as highlighted in yellow in Figure 7, each nonzero element is assigned to exactly one line segment. That is, ALTO imposes a partitioning on a given linearized tensor that generates a disjoint set of non-overlapping and balanced line segments, but it does not guarantee that such a partitioning will divide the multi-dimensional space of the tensor into non-overlapping subspaces. In contrast, the prior compressed formats split the multi-dimensional space into non-overlapping (yet highly imbalanced) regions, namely, tensor slices and fibers in CSF-based formats and multi-dimensional blocks in block-based formats (e.g., HiCOO).
Formally, a set of line segments partitions an ALTO tensor , which encodes -dimensional sparse data, such that and , where . Each line segment l is an ordered set of nonzero elements that are bounded in an -dimensional space by a set of closed mode intervals , where each mode interval is delineated by a start and an end . The intersection of two mode interval sets represents the subspace overlap between their corresponding line segments (partitions), as highlighted in yellow in Figure 7. This overlap information is used to more efficiently resolve conflicts between partitions, as described in 4.2.
@@ 4.2 Adaptive Conflict Resolution
Because processing the nonzero elements of a tensor in parallel (e.g., line 1 in Algorithm 3) can result in write conflicts across threads (e.g., line 3 in Algorithm 3), we devise an adaptive parallel algorithm to handle these conflicts by exploiting the inherent data reuse of target tensors. Our adaptive algorithm chooses between 1) recursively traversing the tensor to maximize the reuse of both the input and output fibers, at the cost of global parallel reduction, or 2) traversing the tensor elements in output-oriented ordering, and only synchronizing at partition boundaries.
Algorithm 4 describes the proposed parallel execution scheme using a representative MTTKRP operation that works on a sparse tensor stored in the ALTO format. After ALTO imposes a partitioning on a sparse tensor, as detailed in 4.1, each partition can be assigned to a different thread. To resolve the update/write conflicts that may happen during parallel sparse tensor computations, ALTO uses an adaptive conflict resolution approach that automatically selects the appropriate tensor traversal and synchronization technique (highlighted by the different gray backgrounds) based on the reuse of the target fibers. This metric represents the average number of nonzero elements per fiber (i.e., the generalization of a matrix row or column) and it is computed in constant time by simply dividing the total number of nonzero elements by the number of fibers along the target mode.
@@
ALTO considers the fiber reuse large enough to use local staging memory for conflict resolution, if the average number of nonzero elements per fiber is more than the maximum cost of using this two-stage (buffered) accumulation process, which consists of initialization (omitted for brevity), local accumulation (line 6 in Algorithm 4), and global accumulation (lines 14–18). In the worst (no reuse) case, the buffered accumulation cost is four memory operations (two read and two write operations). As explained in 4.1, each line segment l is bounded in an -dimensional space by a set of closed mode intervals , which is computed during the partitioning of an ALTO tensor; thus, the size of the temporary storage accessed during the accumulation of l’s updates along a mode is directly determined by the mode interval (see lines 6 and 15).
@@ 4.3 Adaptive Memory Management
In many tensor decomposition algorithms, such as CP-APR, the intermediate results of tensor kernels are typically stored and then reused during the iterative optimization loop (as shown in Algorithm 2). However, storing these important calculations can substantially increase memory traffic and require prohibitive amount of memory, especially for large tensors and high decomposition ranks. Hence, in contrast to the traditional algorithm that pre-computes and reuses the intermediate values (ALTO-PRE), we introduce an ALTO-based algorithm variant that recomputes these values on-the-fly (ALTO-OTF). Moreover, we propose a heuristic to dynamically decide whether to reuse or recompute the intermediate results of tensor kernels based on the characteristics of the target data sets and tensor computations. It is important to note that such pre-computation is different from prior memoization approaches [37, 38], which use a non-trivial decision making process to reduce computations by reformulating tensor operations and reusing intermediate results across modes. In contrast, ALTO-PRE uses easily calculable metrics to decide whether or not to reuse intermediate results within a mode, and it performs the same tensor operations as ALTO-OTF.
To demonstrate our adaptive memory management technique, Algorithm 5 shows how the model update () kernel in CP-APR (Line 10 from Algorithm 2) is parallelized using the ALTO format. The matrix from Line 8 in Algorithm 2 calculates a dense matrix for a given mode that is the Khatri-Rao product (KRP) between all factor matrices excluding the mode- factor matrix. However, not every row of is required for a sparse tensor but only rows that correspond to nonzero elements are necessary and actually calculated, leading to a matrix, where and are the number of nonzero elements and the decomposition rank, respectively. In Algorithm 5, if we select to use pre-computed , the kernel reads in the matrix row corresponding to the nonzero element from memory (Line 7); otherwise, it computes the required KRP from the factor matrices (Line 9) using the delineariezd coordinates. Pre-computing the matrix is simple; in line 8 of Algorithm 2 each nonzero element can calculate its Khatri-Rao product in parallel, using the equation from Line 9 in Algorithm 5.
Next, for each nonzero, the corresponding KRP row is used to update the matrix, and the conflicting updates are resolved using our adaptive conflict resolution as detailed in 4.2. Specifically, if there is high fiber reuse, recursive tensor traversal is conducted and the update is made to the temporary scratchpad memory Temp (Line 11), which is later reduced (Lines 20 to 24) to decrease memory contention; otherwise, output-oriented traversal is used to avoid synchronization and atomic operations are utilized to update the matrix (Line 13) only at the boundaries between different ALTO partitions/threads.
@@ 5 Evaluation
We evaluate ALTO-based tensor algorithms against the state-of-the-art sparse tensor libraries and representations in terms of parallel performance, tensor storage, and format generation cost. We conduct a thorough study of key tensor decomposition operations (2.2) and demonstrate the performance characteristics across the third and fourth generation of Intel Xeon Scalable processors, codenamed Ice Lake (ICX) and Sapphire Rapids (SPR), respectively.
@@ 5.1.1 Platform
The experiments were conducted on an Intel Xeon Platinum 8360Y CPU with Ice Lake (ICX) micro-architecture, and an Intel Xeon Platinum 8470 CPU with Sapphire Rapids (SPR) micro-architecture. The ICX system has 256 GiB main memory and it consists of two sockets, each with a 54 MiB L3 cache and 36 physical cores running at a fixed frequency of 2.4 GHz for accurate measurements. The SPR system also comprises two sockets with 52 physical cores each, and its frequency was fixed to 2.4 GHz. All cores within a socket share a 105 MiB L3 cache and the overall main memory in the SPR server is 1 TB. The experiments use all hardware threads ( and , respectively) on the target platforms. Both servers run AlmaLinux 8.8 Linux distribution and they are configured to enable “Transparent Huge Pages” and to support two and four NUMA domains per socket on ICX and SPR, respectively.
The code is built using Intel C/C++ compiler (v2021.6.0) with the optimization flags -O3 -qopt-zmm-usage=high -xHost to fully utilize vector units. For performance counter measurements and thread pinning, we use the LIKWID tool suite v5.3 [44].
Tensor | Dimensions | #NNZs | Density | Fib. reuse |
lbnl | Limited | |||
nips | High | |||
uber | High | |||
chicago | High | |||
vast | High | |||
darpa | Limited | |||
enron | High | |||
lanl-2 | High | |||
nell-2 | High | |||
fb-m | Limited | |||
flickr | Limited | |||
deli | Medium | |||
nell-1 | Medium | |||
amazon | High | |||
patents | High | |||
High |
@@ 5.1.3 Configurations
We evaluate the proposed ALTO format222Available at: https://github.com/IntelLabs/ALTO, compared to the mode-agnostic COO and HiCOO formats [15, 22] as well as the mode-specific CSF formats [25, 20, 38]. Specifically, we use the latest code of the state-of-the-art sparse tensor libraries for CPUs, namely, ParTI333Available at: https://github.com/hpcgarage/ParTI, SPLATT444Available at: https://github.com/ShadenSmith/splatt, and STeF555Available at: https://github.com/HPCRL/STeF for normally distributed data and SparTen666Available at: https://github.com/sandialabs/sparten for non-negative count data. On the ICX and SPR systems, we evaluate the target data sets that can fit in memory for all tensor libraries. We report the best-achieved performance across the different configurations of the COO format; that is, with or without thread privatization (which keeps local copies of the output factor matrix). For the HiCOO format, its performance and storage are highly sensitive to the block and superblock (SB) sizes, which benefit from tuning. Since the current HiCOO implementation does not auto-tune these parameters, we use a block size of 128 () and two superblock sizes of (“HiCOO-SB10”) and (“HiCOO-SB14”) according to prior work [16]. We evaluate two variants of the mode-specific formats: CSF and CSF with tensor tilling (“CSF-tile”), both of which use representations for an order- sparse tensor to achieve the best performance. For STeF, we use its data movement model to decide which results to memoize and reuse between modes based on the cache size. We report the best performance of STeF across the different cache configurations, which was achieved when setting the cache size to the size of L3 cache on the target ICX and SPR CPUs.
Similar to previous studies [25, 20, 17], the experiments use double-precision arithmetic and -bit (native word) integers. To compute the CP-APR model for non-negative count data, we use -bit integers to represent the input tensor values. While the target data sets require a linearized index of size between and bits, we configured ALTO to select the size of its linearized index to be multiples of the native word size (i.e., and bits) for simplicity. We use a decomposition rank for all experiments and set the maximum number of inner iterations () in CP-APR to , as per prior work [14].